# Solar carbon abundance from 3D non-LTE modelling of the diagnostic lines of the CH molecule

### v1

@@ Abstract
Methods. We use the updated NLTE model molecule from PopaHoppe2023 and different solar 3D radiation-hydrodynamics model atmospheres. The models are contrasted against new spatially-resolved optical solar spectra, and the center-to-limb variation (CLV) of CH lines is studied.
@@
pubyear: 2025pagerange: Solar carbon abundance from 3D non-LTE modelling of the diagnostic lines of the CH moleculeA.5
@@ 1 Introduction
The CH molecule, with its rich spectral features, is an invaluable tool for tracing carbon (C) abundances in stellar atmospheres. Starting from early studies like Russell1934, the CH spectral lines, especially the critical G-band of CH at $4300\text{\,}\mathrm{\text{Å}}$, were used to determine C abundances in metal-poor stars, red giants, stellar populations (ShetroneSmith1999; YoonBeers2016; YoonBeers2018; LucatelloBeers2006; StarkenburgHill2013; PlaccoKennedy2011; HansenNordstrom2016; Suarez-AndresIsraelian2017).
The G-band is particularly important in spectra of metal-poor stars, where atomic carbon (C I) lines are weak and difficult to disentangle from other features (e.g. WheelerSneden1989; AlexeevaMashonkina2015). The solar C abundance is important in the calculation of Standard Solar Models (SSM), which describe the structure of the solar interior and its evolution (e.g. SerenelliBasu2009; PinsonneaultDelahaye2009; VinyolesSerenelli2017).
@@
Also for the Sun, detailed spectral analysis of CH lines in 1D LTE and 3D LTE was performed. The 1D LTE studies include GrevesseLambert1991; AlexeevaMashonkina2015, whereas 3D LTE studies are AsplundGrevesse2005; AsplundGrevesse2009; AsplundAmarsi2021. The 3D LTE studies found qualitatively different effects of 3D convection in the analysis of the diagnostic CH lines in the optical and infra-red wavelength ranges. Whereas (AsplundGrevesse2005) found that 3D LTE value of A(C) is within 0.01 dex of the 1D LTE result, the most recent analysis AsplundAmarsi2021 found a higher 3D LTE A(C) compared to the value obtained with a 1D MARCS models by 0.03-0.07 dex. However, the effects of departures from LTE have not been explored in 3D in any detailed systematic approach to date.
@@
In the NLTE framework, interactions between the radiation field and the gas in stellar atmospheres are calculated by solving the coupled equations of radiation transfer and statistical equilibrium using realistic cross-sections and reaction rates from the lab or theoretical studies (e.g. Asplund2005; HubenyMihalas2014; BergemannNordlander2014; LindAmarsi2024). These are vastly more complex compared to the very simple LTE approximation, which assumes the state of maximum entropy. Studies investigating NLTE effects in molecules have been scarce and focussed on CO (AyresWiedemann1989; SchweitzerHauschildt2000; BerknerSchweitzer2015) and H2O (LambertJosselin2013), but also with highly simplified molecular models, ignoring detailed key reactions like state-resolved photo-dissociation cross-sections or charge exchange reactions. Recently, NLTE effects in the CH diagnostic spectral lines were studied with 1D stellar atmosphere models in our recent work PopaHoppe2023 (hereafter PH23).
@@
In this paper, we provide a systematic analysis of the solar C abundance based on molecular CH lines in the solar spectrum. Different from the previous work, we use NLTE modelling with 3D RHD simulations of solar convection, and we contrast the predictions of models against new spatially-resolved observations of the optical wavelength range (EllwarthSchafer2023; ReinersYan2023). These center-to-limb spectra ($4200\text{\,}$ to $8000\text{\,}\mathrm{\text{Å}}$) cover the optical range of the CH G-band, allowing an entirely novel test of the solar spectroscopic models. We explore the consistency of C abundances retrieved from spectra taken at different angles, and explore the physics underlying 1D LTE, 1D NLTE, 3D LTE and 3D NLTE line formation of the diagnostic CH lines. Finally, we provide the solar 3D NLTE C abundances based on optical and IR lines, and analyse the new value in the context of other measurements and probes, including the solar wind (vonSteigerZurbuchen2016) and data obtained from the Borexino particle physics experiment (Gonzalez-GarciaMaltoni2024).
@@
Figure 1: A view of the solar disk by the Helioseismic and Magnetic Imager (image by NASA’s Scientific Visualization Studio, nasa_sun_image). Overlayed are concentric rings showing viewing positions of constant $\mu$-angle, while the green scale shows corresponding relative radii. We also show the finite fiber size of $32.5\text{\,}\arcsec$ of the IAG instrument (SchaferRoyen2020).
@@
Figure 2: Comparison of synthetic and observed solar G-band intensity at wavelength $4306.8\pm$$15\text{\,}\mathrm{\text{Å}}$ at different positions across the solar disc (different rows). The synthetic G-band has been convolved with the SST photometric filter transmission function (CarlssonStein2004). The G-band observations were acquired with the Swedish 1-m Solar Telescope (SST, Scharmer2003a) on La Palma in 2003-2004 (Berger2004; Lin2005, Van Der Voort priv. comm.). For more detailed information on these observations we refer the reader to https://www.isf.astro.su.se/data1/gallery/. We show the disc center $\mu=1$ at the bottom and two angles approaching the limb $\mu=0.92,\mu=0.55$ above. The black and white levels of the observations were adjusted to account for the different processing of the individual measurements.
@@ 2.1 Observational data
The observed spatially-resolved spectroscopic data around $4300\text{\,}\mathrm{\text{Å}}$ are taken from the solar intensity FTS atlas by EllwarthSchafer2023 taken at the Institute for Astrophysics and Geophysics(IAG), Göttingen. The atlas covers the full wavelength range between $4200\text{\,}\mathrm{-}$ $8000\text{\,}\mathrm{\text{Å}}$ and contains observations of the solar disk at 14 different viewing angles between $\mu=1.0$ and $\mu=0.2$. The $\mu$-angle and relative radius $r/r_{\odot}$ are defined as
@@
For the infrared transitions at (32 94632\,946-$37\,955\text{\,}\mathrm{\text{Å}}$) we use the Spacelab-3 ATMOS atlas ($\mu=0.935$) (FarmerNorton1989) as uploaded by SeoKim2007 (file solspec2b_1.txt)  [^1] https://www.sciencedirect.com/science/article/pii/S0019103507003089?via%3Dihub#ec005 as NASA’s original links are deprecated. Due to wiggles in the continuum we renormalise the atlas utilising the neural network normalisation tool SUPPNet (RozanskiNiemczura2022) with the sampling set to $0.05\text{\,}\mathrm{\text{Å}}$, a smoothing factor of $2$ and no further manual adjustment of the knots of the continuum spline.
The G-band observations in Fig. 2 were acquired with the Swedish 1-m Solar Telescope (SST, Scharmer2003a) on La Palma in 2003-2004. The images were acquired with an interference filter centred on $4305\text{\,}\mathrm{\text{Å}}$ and with bandwidth of 11 $1\text{\,}\mathrm{\text{Å}}$. Seeing conditions were such that near-diffraction limited imaging was achieved with the aid of adaptive optics (Scharmer2003b) and image restoration (Löfdahl2002; VanNoort2005). The $\mu=1$ and the $\mu=0.92$ observations were published in Berger2004 and Lin2005.
@@ 2.2 The Code
We use our updated version of the 3D NLTE radiative transfer code MULTI3D within the DISPATCH framework, which provides a task-based parallelisation scheme through MPI as well as openMP (NordlundRamsey2018).
Even though the original version of MULTI3D (LeenaartsCarlsson2009) already featured spatial and frequency domain decomposition, the task-based approach of the DISPATCH framework has several advantages which will be discussed in an accompanying paper (Hoppe in prep.). Besides the new parallelisation scheme we have updated the background bound-free opacities (see EitnerBergemann2024). We are confident that the radiative transfer, EOS, background opacities and broadening recipes work reliably, as we tested synthetic CH line profiles computed with MULTI3D against the independent spectrum synthesis code TURBOSPECTRUM in 1D LTE (see Fig. 12 in Appendix).
In the NLTE calculations for the CH molecule, we included background bound-bound opacities during the statistical equilibrium calculations via interpolation of opacity tables precomputed from MARCS (GustafssonEdvardsson2008).
For the purposes of this work, we employ auto-preconditioning (RybickiHummer1992, Section 2.4), which suppresses only the feedback in the radiative rates of each line with itself. This approach is also used in other 3D NLTE codes such as Balder (AmarsiBarklem2019). In our case, it is adopted to reduce the computational overhead, which becomes significant when dealing with a large number of energy levels in the full-preconditioning scheme (RybickiHummer1992, Section 2.3) originally implemented in MULTI3D.
@@ 2.3 Diagnostic CH lines used for the solar C abundance analysis
Our line selection is based on the one by AsplundGrevesse2005 (AGS05) who focused on 9 optical A-X transitions and 102 infrared rotation-vibration transitions to determine the solar C abundance. While we adopted the same 9 optical transitions (table 1), we reduced the infrared line sample down to 50 lines (table 3). We removed lines from the infrared sample where the continuum normalisation by SUPPNet was unsatisfactory or the line was positioned in the wing of a stronger line after visual inspection. We also removed those lines resulting from superposition of multiple CH transitions being spaced less than $0.7\text{\,}\mathrm{\text{Å}}$ apart. This was done because the theoretical line centers from the MasseronPlez2014 line list do not perfectly match the observed line centers to the $0.1\text{\,}\mathrm{\text{Å}}$ level. For single lines this can be accounted for by slightly shifting the line in the fit. However, when the spacing between the centers of two blended lines is off, the compound shape of the lines does not necessarily fit the observed shape. All diagnostic transitions are highlighted in the Grotrian diagram of the CH molecule in Fig. 3 showing its different electronic, vibrational and rotational enery states.
@@ 2.4.1 Superlevels
In order to perform 3D NLTE calculations, the model molecule had to be reduced, which is a common procedure for atoms (BergemannLind2012). This is because it is in practice extremely time-consuming to do matrix inversions for a over a few hundred energy states coupled by radiation and collisional transitions.
Therefore, we combined energy levels that are very close in excitation energy value into superlevels (e.g., HubenyMihalas2014). For this we used a K-means clustering algorithm in Python to group and merge the levels with similar excitation energy. Besides merging energy levels we also took great care to merge photo-dissociation transitions accordingly. After reducing the CH molecule to $294$ energy levels the computational cost of the matrix inversions was negligible, while the non-LTE effects on the populations were preserved, as discussed in section A.2 in the Appendix.
@@ 2.4.3 Collisional transitions
Theoretical excitation rates have been presented for CH collisions with He (MarinakisDean2015) as well as for collisions with H and H2 (Dagdigian2016; Dagdigian2017; Dagdigian2018). These studies have probed the temperature range between 10-$300\text{\,}\mathrm{K}$ relevant for excitation of CH molecules in the interstellar medium. The behaviour of the collision rates as a function of temperature differs drastically between the different colliders. While collisions with H and He show a strong temperature dependence, rate constants for collisions with H2 are approximately constant with temperature. The extrapolation to temperatures in stellar atmospheres has been discouraged by the authors (see PopaHoppe2023). More recently, empirical recipes for collisions between CH and He as well as H2 have been presented by YanWang2025 based on a compilation of experimental collision data for NO, OH and CH molecules with a range of different colliders. This study focuses on the forward modelling of laser-induced-fluorescence (LIF) which aids the analysis of experimental LIF measurements, for example in combustion systems. While their recipes are applicable in a temperature range of 200-$3000\text{\,}\mathrm{K}$ , many of the recipe parameters for CH have been deduced by analogy from NO or OH.
@@ 2.5 Model atmospheres
We compare the results from hydrostatic and 3D RHD atmospheric models of the Sun. We employ 11 snapshots from a solar M3DIS simulation (EitnerBergemann2024) and 13 snapshots from a solar STAGGER simulation (MagicCollet2013), both of which are 3D RHD simulations. It has been demonstrated that sampling ten statistically independent snapshots is sufficient to achieve accurate line-profile fits, with a maximum flux error of 0.005 (RodriguezDiazLagae2024). Earlier solar calculations already showed that using as few as five snapshots yields reliable results (BergemannGallagher2019; GallagherBergemann2020). For comparison, we use a 1D hydrostatic MARCS solar model (GustafssonEdvardsson2008) with a microturbulence of $v_{\rm mic}=$1\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$. Additionally, we use two average 3D (or <3D>) models, which are 1D models derived by averaging 3D hydrodynamical snapshots over surfaces of equal optical depth at $5000\text{\,}\mathrm{\text{Å}}$. These are the <M3DIS > and <CO5BOLD > models. When using the <3D> atmospheres, we also assume a $v_{\rm mic}=$1\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$. We did not consider the CO5BOLD model in full 3D, as only the <3D> version of the model is available to us (MaggBergemann2022). For this study, we resample both hydrostatic and hydrodynamic atmospheres based on their electron pressure gradient to ensure detailed sampling in the photosphere. This resampling has negligible impact on the spectra of 1D models, which are usually tabulated in equal optical depth steps, but is common practice in the case of hydrodynamical atmospheres which have to extend to optically thick layers $\log\tau\gg 1$ at the expense of fine sampling in the photosphere (AmarsiNordlander2018; RodriguezDiazLagae2024). The temperature structure of the M3DIS simulation is presented in Fig. 4. Average temperature differences between the rest of the employed atmospheres and how they affect the CH abundance are shown in Fig. 13 in the Appendix.

### v2

@@ Abstract
Methods. We use the updated NLTE model molecule from Popa et al. (2023) and different solar 3D radiation-hydrodynamics model atmospheres. The models are contrasted against new spatially-resolved optical solar spectra, and the center-to-limb variation (CLV) of CH lines is studied.
@@
pubyear: 2026pagerange: Solar carbon abundance from 3D non-LTE modelling of the diagnostic lines of the CH moleculeA.5
@@ 1 Introduction
The CH molecule, with its rich spectral features, is an invaluable tool for tracing carbon (C) abundances in stellar atmospheres. Starting from early studies like Russell (1934), the CH spectral lines, especially the critical G-band of CH at $4300\text{\,}\mathrm{\text{Å}}$, were used to determine C abundances in metal-poor stars, red giants, stellar populations (Shetrone et al., 1999; Yoon et al., 2016, 2018; Lucatello et al., 2006; Starkenburg et al., 2013; Placco et al., 2011; Hansen et al., 2016; Suárez-Andrés et al., 2017).
The G-band is particularly important in spectra of metal-poor stars, where atomic carbon (C I) lines are weak and difficult to disentangle from other features (e.g. Wheeler et al., 1989; Alexeeva and Mashonkina, 2015). The solar C abundance is important in the calculation of Standard Solar Models (SSM), which describe the structure of the solar interior and its evolution (e.g. Serenelli et al., 2009; Pinsonneault and Delahaye, 2009; Vinyoles et al., 2017).
@@
Also for the Sun, detailed spectral analysis of CH lines in 1D LTE and 3D LTE was performed. The 1D LTE studies include Grevesse et al. (1991); Alexeeva and Mashonkina (2015), whereas 3D LTE studies are Asplund et al. (2005, 2009, 2021). The 3D LTE studies found qualitatively different effects of 3D convection in the analysis of the diagnostic CH lines in the optical and infra-red wavelength ranges. Whereas (Asplund et al., 2005) found that the 3D LTE value of A(C) is within 0.01 dex of the 1D LTE result, the most recent analysis Asplund et al. (2021) found a higher 3D LTE A(C) compared to the value obtained with a 1D MARCS models by 0.03-0.07 dex. However, the effects of departures from LTE have not been explored in 3D in any detailed systematic approach to date.
@@
In the NLTE framework, interactions between the radiation field and the gas in stellar atmospheres are calculated by solving the coupled equations of radiation transfer and statistical equilibrium using realistic cross-sections and reaction rates from the lab or theoretical studies (e.g. Asplund, 2005; Hubeny and Mihalas, 2014; Bergemann and Nordlander, 2014; Lind and Amarsi, 2024). These are vastly more complex compared to the very simple LTE approximation, which assumes the state of maximum entropy. Studies investigating NLTE effects in molecules have been scarce and focussed on CO (Ayres and Wiedemann, 1989; Schweitzer et al., 2000; Berkner et al., 2015) and H2O (Lambert et al., 2013), but also with highly simplified molecular models, ignoring detailed key reactions like state-resolved photo-dissociation cross-sections or charge exchange reactions. Recently, NLTE effects in the CH diagnostic spectral lines were studied with 1D stellar atmosphere models in our recent work Popa et al. (2023) (hereafter PH23).
@@
In this paper, we provide a systematic analysis of the solar C abundance based on molecular CH lines in the solar spectrum. Different from the previous work, we use NLTE modelling with 3D RHD simulations of solar convection, and we contrast the predictions of models against new spatially-resolved observations of the optical wavelength range (Ellwarth et al., 2023; Reiners et al., 2023). These center-to-limb spectra ($4200\text{\,}$ to $8000\text{\,}\mathrm{\text{Å}}$) cover the optical range of the CH G-band, allowing an entirely novel test of the solar spectroscopic models. We explore the consistency of C abundances retrieved from spectra taken at different angles, and explore the physics underlying 1D LTE, 1D NLTE, 3D LTE and 3D NLTE line formation of the diagnostic CH lines. Finally, we provide the solar 3D NLTE C abundances based on optical and IR lines, and analyse the new value in the context of other measurements and probes, including the solar wind (von Steiger and Zurbuchen, 2016) and data obtained from the Borexino particle physics experiment (Gonzalez-Garcia et al., 2024).
@@
Figure 1: A view of the solar disk by the Helioseismic and Magnetic Imager (image by NASA’s Scientific Visualization Studio, Bridgman et al., 2021). Overlayed are concentric rings showing viewing positions of constant $\mu$-angle, while the green scale shows corresponding relative radii. We also show the finite fiber size of $32.5\text{\,}\arcsec$ of the IAG instrument (Schäfer et al., 2020).
@@
Figure 2: Comparison of synthetic and observed solar G-band intensity at wavelength $4306.8\pm$$15\text{\,}\mathrm{\text{Å}}$ at different positions across the solar disc (different rows). The synthetic G-band has been convolved with the SST photometric filter transmission function (Carlsson et al., 2004). The G-band observations were acquired with the Swedish 1-m Solar Telescope (SST, Scharmer et al. (2003a)) on La Palma in 2003-2004 (Berger et al., 2004; Lin et al., 2005, Van Der Voort priv. comm.). For more detailed information on these observations we refer the reader to https://www.isf.astro.su.se/data1/gallery/. We show the disc center $\mu=1$ at the bottom and two angles approaching the limb $\mu=0.92,\mu=0.55$ above. The black and white levels of the observations were adjusted to account for the different processing of the individual measurements.
@@ 2.1 Observational data
The observed spatially-resolved spectroscopic data around $4300\text{\,}\mathrm{\text{Å}}$ are taken from the solar intensity FTS atlas by Ellwarth et al. (2023) taken at the Institute for Astrophysics and Geophysics(IAG), Göttingen. The atlas covers the full wavelength range between $4200\text{\,}\mathrm{-}$ $8000\text{\,}\mathrm{\text{Å}}$ and contains observations of the solar disk at 14 different viewing angles between $\mu=1.0$ and $\mu=0.2$. The $\mu$-angle and relative radius $r/r_{\odot}$ are defined as
@@
For the infrared transitions at (32 94632\,946-$37\,955\text{\,}\mathrm{\text{Å}}$) we use the Spacelab-3 ATMOS atlas ($\mu=0.935$) (Farmer and Norton, 1989) as uploaded by Seo et al. (2007) (file solspec2b_1.txt)  [^1] https://www.sciencedirect.com/science/article/pii/S0019103507003089?via%3Dihub#ec005 as NASA’s original links are deprecated. Due to wiggles in the continuum we renormalise the atlas utilising the neural network normalisation tool SUPPNet (Różański et al., 2022) with the sampling set to $0.05\text{\,}\mathrm{\text{Å}}$, a smoothing factor of $2$ and no further manual adjustment of the knots of the continuum spline.
The G-band observations in Fig. 2 were acquired with the Swedish 1-m Solar Telescope (SST, Scharmer et al. (2003a)) on La Palma in 2003-2004. The images were acquired with an interference filter centred on $4305\text{\,}\mathrm{\text{Å}}$ and with bandwidth of 11 $1\text{\,}\mathrm{\text{Å}}$. Seeing conditions were such that near-diffraction limited imaging was achieved with the aid of adaptive optics (Scharmer et al., 2003b) and image restoration (Löfdahl2002; Van Noort et al., 2005). The $\mu=1$ and the $\mu=0.92$ observations were published in Berger et al. (2004) and Lin et al. (2005).
@@ 2.2 The Code
We use our updated version of the 3D NLTE radiative transfer code MULTI3D within the DISPATCH framework, which provides a task-based parallelisation scheme through MPI as well as openMP (Nordlund et al., 2018).
Even though the original version of MULTI3D (Leenaarts and Carlsson, 2009) already featured spatial and frequency domain decomposition, the task-based approach of the DISPATCH framework has several advantages which will be discussed in an accompanying paper (Hoppe in prep.). Besides the new parallelisation scheme we have updated the background bound-free opacities (see Eitner et al., 2024). We are confident that the radiative transfer, EOS, background opacities and broadening recipes work reliably, as we tested synthetic CH line profiles computed with MULTI3D against the independent spectrum synthesis code TURBOSPECTRUM in 1D LTE (see Fig. 12 in Appendix).
In the NLTE calculations for the CH molecule, we included background bound-bound opacities during the statistical equilibrium calculations via interpolation of opacity tables precomputed from MARCS (Gustafsson et al., 2008).
For the purposes of this work, we employ auto-preconditioning (Rybicki and Hummer, 1992, Section 2.4), which suppresses only the feedback in the radiative rates of each line with itself. This approach is also used in other 3D NLTE codes such as Balder (Amarsi et al., 2019). In our case, it is adopted to reduce the computational overhead, which becomes significant when dealing with a large number of energy levels in the full-preconditioning scheme (Rybicki and Hummer, 1992, Section 2.3) originally implemented in MULTI3D.
@@ 2.3 Diagnostic CH lines used for the solar C abundance analysis
Our line selection is based on the one by Asplund et al. (2005) (AGS05) who focused on 9 optical A-X transitions and 102 infrared rotation-vibration transitions to determine the solar C abundance. While we adopted the same 9 optical transitions (table 1), we reduced the infrared line sample down to 50 lines (table 3). We removed lines from the infrared sample where the continuum normalisation by SUPPNet was unsatisfactory or the line was positioned in the wing of a stronger line after visual inspection. We also removed those lines resulting from superposition of multiple CH transitions being spaced less than $0.7\text{\,}\mathrm{\text{Å}}$ apart. This was done because the theoretical line centers from the Masseron et al. (2014) line list do not perfectly match the observed line centers to the $0.1\text{\,}\mathrm{\text{Å}}$ level. For single lines this can be accounted for by slightly shifting the line in the fit. However, when the spacing between the centers of two blended lines is off, the compound shape of the lines does not necessarily fit the observed shape. All diagnostic transitions are highlighted in the Grotrian diagram of the CH molecule in Fig. 3 showing its different electronic, vibrational and rotational enery states.
@@ 2.4.1 Superlevels
In order to perform 3D NLTE calculations, the model molecule had to be reduced, which is a common procedure for atoms (Bergemann et al., 2012). This is because it is in practice extremely time-consuming to do matrix inversions for a over a few hundred energy states coupled by radiation and collisional transitions.
Therefore, we combined energy levels that are very close in excitation energy value into superlevels (e.g., Hubeny and Mihalas, 2014). For this we used a K-means clustering algorithm in Python to group and merge the levels with similar excitation energy. Besides merging energy levels we also took great care to merge photo-dissociation transitions accordingly. After reducing the CH molecule to $294$ energy levels the computational cost of the matrix inversions was negligible, while the non-LTE effects on the populations were preserved, as discussed in section A.2 in the Appendix.
@@ 2.4.3 Collisional transitions
Theoretical excitation rates have been presented for CH collisions with He (Marinakis et al., 2015) as well as for collisions with H and H2 (Dagdigian, 2016, 2017, 2018). These studies have probed the temperature range between 10-$300\text{\,}\mathrm{K}$ relevant for excitation of CH molecules in the interstellar medium. The behaviour of the collision rates as a function of temperature differs drastically between the different colliders. While collisions with H and He show a strong temperature dependence, rate constants for collisions with H2 are approximately constant with temperature. The extrapolation to temperatures in stellar atmospheres has been discouraged by the authors (see Popa et al., 2023). More recently, empirical recipes for collisions between CH and He as well as H2 have been presented by Yan and Wang (2025) based on a compilation of experimental collision data for NO, OH and CH molecules with a range of different colliders. This study focuses on the forward modelling of laser-induced-fluorescence (LIF) which aids the analysis of experimental LIF measurements, for example in combustion systems. While their recipes are applicable in a temperature range of 200-$3000\text{\,}\mathrm{K}$ , many of the recipe parameters for CH have been deduced by analogy from NO or OH.
@@ 2.5 Model atmospheres
We compare the results from hydrostatic and 3D RHD atmospheric models of the Sun. We employ 11 snapshots from a solar M3DIS simulation (Eitner et al., 2024) and 13 snapshots from a solar STAGGER simulation (Magic et al., 2013), both of which are 3D RHD simulations. It has been demonstrated that sampling ten statistically independent snapshots is sufficient to achieve accurate line-profile fits, with a maximum flux error of 0.005 (Rodríguez Díaz et al., 2024). Earlier solar calculations already showed that using as few as five snapshots yields reliable results (Bergemann et al., 2019; Gallagher et al., 2020). For comparison, we use a 1D hydrostatic MARCS solar model (Gustafsson et al., 2008) with a microturbulence of $v_{\rm mic}=$1\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$. Additionally, we use two average 3D (or <3D>) models, which are 1D models derived by averaging 3D hydrodynamical snapshots over surfaces of equal optical depth at $5000\text{\,}\mathrm{\text{Å}}$. These are the <M3DIS > and <CO5BOLD > models. When using the <3D> atmospheres, we also assume a $v_{\rm mic}=$1\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$. We did not consider the CO5BOLD model in full 3D, as only the <3D> version of the model is available to us (Magg et al., 2022). For this study, we resample both hydrostatic and hydrodynamic atmospheres based on their electron pressure gradient to ensure detailed sampling in the photosphere. This resampling has negligible impact on the spectra of 1D models, which are usually tabulated in equal optical depth steps, but is common practice in the case of hydrodynamical atmospheres which have to extend to optically thick layers $\log\tau\gg 1$ at the expense of fine sampling in the photosphere (Amarsi et al., 2018; Rodríguez Díaz et al., 2024). The temperature structure of the M3DIS simulation is presented in Fig. 4. Average temperature differences between the rest of the employed atmospheres and how they affect the CH abundance are shown in Fig. 13 in the Appendix.